OpenStudy (anonymous):

Virago bought 78 pots for $312. Write the two rates (ratios) implied by this statement. What would she have had to pay for 400 pots?

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@dilphill31

OpenStudy (igreen):

Just divide.. 312 / 78 = ?

OpenStudy (igreen):

And 78 / 312

OpenStudy (anonymous):

ok one sec

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

and 0.25

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

Find the rate per pot: write the ratio $312/78 pots. We want to know the dollar amount for 400 pots. Write another ratio, using x for the dollar amount, over 400: $312/78 pots = x/ 400 pots.

OpenStudy (mathmale):

Solve this equation of ratios for x. Be certain to check y our work.

OpenStudy (anonymous):

well I divided 78/$312 and I got 0.25 and I also divided $312/78 and got 4

OpenStudy (mathmale):

What was your objective in doing that?

OpenStudy (anonymous):

the i green guy told me to do that

OpenStudy (mathmale):

Of great interest is the unit cost per pot. You obtain that as follows:\[\frac{ $312 }{ 78 pots }\]

OpenStudy (mathmale):

Obviously, 400 pots are going to cost more than 78 pots. Write another ratio: \[\frac{ $x }{ 400 }\]

OpenStudy (mathmale):

Set those 2 ratios equal to each other; then cross-multiply. Are you familiar with cross mult?

OpenStudy (anonymous):

yes I am. One second.

OpenStudy (mathmale):

\[\frac{ $312 }{ 78 }=\frac{ x$ }{ 400 }\]

OpenStudy (mathmale):

Mult x by 78. Mult $312 by 400. Set these 2 products = to each other. Solve the resulting eq'n for x.

OpenStudy (anonymous):

Okay, thank you!

OpenStudy (mathmale):

You're welcome, Ricky! Take care.